### v4

# Planning-Aware Code Infilling via Horizon-Length Prediction

Abstract
Fill-in-the-Middle (FIM), or infilling, has become integral to code language models, enabling generation of missing code given both left and right contexts. However, the current FIM training paradigm which performs next-token prediction (NTP) over reordered sequence often leads to models struggling to generate content that aligns well with the surrounding context. We hypothesize that NTP alone is insufficient for models to learn effective planning conditioned on the distant right context, a critical factor for successful code infilling. To overcome this, we propose Horizon-Length Prediction (HLP), a novel training objective that teaches models to predict the number of remaining middle tokens at each step. HLP advances FIM with lookahead planning, enabling models to inherently learn infilling boundaries for arbitrary left and right contexts without relying on dataset-specific post-processing. Our evaluation across different model families and sizes shows that HLP significantly improves FIM performance by up to 24% relatively on diverse benchmarks, across file-level and repository-level. Furthermore, the enhanced planning capability gained through HLP boosts model performance on code reasoning. Importantly, HLP incurs negligible training overhead and no additional inference cost, ensuring its practicality for real-world scenarios.
Planning-Aware Code Infilling via Horizon-Length Prediction
Yifeng Ding1††thanks: Work done at AWS AI Labs. Hantian Ding2 Shiqi Wang3∗ Qing Sun2 Varun Kumar2 Zijian Wang3∗ 1University of Illinois Urbana-Champaign 2AWS AI Labs 3Meta yifeng6@illinois.edu {dhantian,qinsun,kuvrun}@amazon.com {tcwangshiqi,zijianwang}@meta.com
1 Introduction
Fill-in-the-Middle (FIM), or infilling, has become essential for modern code development, where programmers frequently need to insert or modify code between existing sections rather than writing linearly from start to end (openaifim; incoder). While large language models have shown remarkable capabilities in code generation, the FIM task poses unique challenges that go beyond traditional left-to-right generation. A model must not only generate code that follows from the preceding context (), but also smoothly connect to the subsequent code () – a task that requires careful planning and foresight.
Current approaches to FIM typically reorder the input sequence and apply standard next-token prediction (NTP) training (starcodertwo; dscoder; dscodervtwo; qwentwofive). However, as illustrated in Figure 1, this methodology has a fundamental limitation: models struggle to maintain coherence over longer sequences and often fail to create smooth transitions to the right context (openaifim). While existing benchmarks attempt to address this through rule-based post-processing (e.g., truncating generated code based on line count (repoeval; repoformer) or program structure (crosscodeeval; safim)), such methods rely on dataset-specific assumptions that do not generalize to real-world scenarios where both left and right contexts can be arbitrary.
| Post-processing Criteria | |
| RepoEval (repoeval) | Truncate generation to the same number of lines as in ground truth. |
| CrossCodeLongEval (repoformer) | |
| CrossCodeEval (crosscodeeval) | Truncate generation at the first complete statement. |
| SAFIM (safim) | Stop when the target program structure in ground truth is generated. |
We hypothesize that the core challenge lies in the model’s limited ability to plan ahead while filling in the middle. Standard next-token prediction training only requires considering one token at a time, but successful code infilling demands awareness of the entire missing section to ensure both local coherence and proper connection to the right context. This planning capability is particularly crucial because generated code must satisfy both syntactic requirements and semantic constraints from both surrounding contexts.
To address this limitation, we propose Horizon-Length Prediction (HLP), a novel training objective that teaches models to predict the number of remaining tokens needed to complete the middle section at each generation step. HLP advances NTP by encouraging models to develop awareness of the generation horizon and plan accordingly. Unlike post-processing approaches, HLP is generalizable as it does not rely on any task-specific knowledge. Instead, it strengthens models’ innate ability to plan and execute coherent code completions.
Our comprehensive evaluation across different models and model sizes demonstrates that HLP significantly improves FIM performance, achieving up to 24% relative improvement on diverse benchmarks at both file-level and repository-level without relying on any dataset-specific post-processing. Furthermore, the enhanced planning capability gained through HLP training also boosts model performance on code reasoning tasks. Importantly, HLP incurs negligible training overhead and no additional inference cost, making it practical for real-world deployment.
Our key contributions are as follows:
We identify planning capability as a fundamental bottleneck in current FIM approaches and quantitatively demonstrate how post-processing methods in existing benchmarks obscure this critical limitation.
We propose Horizon-Length Prediction (HLP), a novel training objective that advances fill-in-the-middle capability by teaching models to plan over arbitrarily long horizons, with negligible training and zero inference overhead.
We demonstrate up to 24% improvement in FIM performance across multiple benchmarks and model families without relying on any post-processing.
We show that HLP’s benefits extend beyond FIM to improve performance on code reasoning tasks, and present various analyses that illuminate the underlying mechanism that enable these improvements.
2 Post-processing for Fill-in-the-Middle
Most existing FIM works rely on post-processing to truncate code completions generated by LLMs for infilling tasks (safim; repoeval; crosscodeeval; repoformer). While such post-processing can enhance the FIM performance, we argue that they fundamentally depend on dataset-specific assumptions that make them impractical for real-world scenarios (§2.1). Through evaluation, we show that the performance of FIM existing code models drops significantly without post-processing, suggesting that post-processing conceals the fundamental struggles of models with code filling (§2.2). Furthermore, we show that this limitation stems from models’ inability to plan coherent completions given a fixed – a challenge that post-processing alone cannot address.
2.1 Post-processing Requires Task-Specific Knowledge
Post-processing methods adopted by recent FIM benchmarks typically assume a certain completion type and perform rule-based truncation accordingly. Table 1 summarizes the post-processing criteria of four popular FIM benchmarks, highlighting the specific rule used for each dataset. These criteria do not transfer across datasets, nor are they generalizable to FIM in the real-world scenario where both left and right contexts can be arbitrary. Given the complexity of programming languages, developing universally applicable post-processing rules for infilling is infeasible. Instead, models need to learn the intrinsic patterns that make for good completions.
| SAFIM | Avg | ||||
| Algo | Algo | Control | API | ||
| DS-1.3B | |||||
| w/ post | 43.9 | 49.2 | 55.6 | 62.9 | 52.9 |
| w/o post | 39.8 | 42.4 | 52.4 | 56.1 | 47.7 |
| rel. diff | -9.3% | -13.8% | -5.8% | -10.8% | -9.9% |
| DS-6.7B | |||||
| w/ post | 54.9 | 58.9 | 68.1 | 71.0 | 63.2 |
| w/o post | 53.4 | 56.7 | 66.6 | 69.0 | 61.4 |
| rel. diff | -2.7% | -3.7% | -2.2% | -2.8% | -2.8% |
| SC2-3B | |||||
| w/ post | 48.1 | 53.5 | 60.1 | 68.4 | 57.5 |
| w/o post | 45.4 | 49.7 | 57.1 | 61.3 | 53.4 |
| rel. diff | -5.6% | -7.1% | -5.0% | -10.4% | -7.2% |
| SC2-7B | |||||
| w/ post | 50.4 | 55.8 | 62.3 | 70.3 | 59.7 |
| w/o post | 48.4 | 53.1 | 60.4 | 63.9 | 56.5 |
| rel. diff | -4.0% | -4.8% | -3.0% | -9.1% | -5.4% |
2.2 LLMs Fail to Plan Coherent Completions
To further demonstrate to what extent post-processing conceals LLMs’ inability of connecting to , we conduct a comprehensive experiment on SAFIM. We compare FIM performance of four different code LLMs, with or without post-processing. As shown in Table 2, removing post-processing leads to up to a substantial 13.8% Pass@1 drop across all models. This reveals that current models have not truly mastered the fundamental task of generating code that properly connects and contexts. Post-processing creates an illusion of competence by artificially “fixing” problematic generations rather than addressing the core limitation.
2.3 FIM Requires Planning Capability
The challenges in FIM stem from models’ inability to plan coherent completions. Consider the example in Figure 1, where a model needs to implement a speech recognition initialization: while the model demonstrates knowledge of the required components (using Recognizer), it fails to order them properly. Without careful planning, it prematurely places the Recognizer call, leading to both syntactic and semantic errors. Importantly, this type of failure cannot be fixed through post-processing, as truncating the generation would lose essential statements while keeping it results in invalid code.
Such an example illustrates that successful FIM requires not only knowledge of individual components but also the model’s ability to plan coherent sequences that smoothly connect to both contexts. The model must reason about the entire completion considering both local coherence and global structure, and that points to a clear need for models to develop intrinsic planning capabilities to succeed at FIM.
3 Horizon-Length Prediction
Given a document that contains tokens , existing FIM training scheme can be divided into three steps: (1) Split the document into three parts: -- 111We opt to use PSM setting in this work given our base models DeepSeek-Coder and StarCoder2 were both pre-trained with PSM. We expect that our method is generalizable to SPM setting as well., (2) Construct a new FIM-style document by reordering the three parts as --, and (3) Conduct next-token prediction (NTP) training on the document .
Specifically, we define the three parts in document as , , and . Then, the new document will be formatted as follows:
where the last step re-indexes the leading tokens up until <mid> to to focus on the FIM part, as LLMs are expected to start infilling after <mid> token and to end generation with <eoi> token to connect to accurately.
Next-token prediction (NTP) training is conducted on the document , which aims to minimize the following cross-entropy loss (where refers to the LLMs being trained):
While NTP provides basic FIM capabilities, it has a fundamental limitation: the model only learns to predict one token at a time without developing awareness of the overall horizon. This makes it difficult to plan coherent sequences that properly connect to both contexts.
Horizon-Length Prediction (HLP). To mitigate this issue, we propose horizon-length prediction (HLP) as an auxiliary training objective. As shown in Figure 2, the key idea is to teach models to predict the number of remaining tokens needed to complete the section at each generation step. This creates an explicit training signal for planning awareness.
Specifically, at each position in , HLP predicts
where is the total length of . This normalized value represents the portion of that remains to be generated and ensures the target is always within interval regardless of the model’s context window size.
HLP is implemented as a linear layer on top of the transformer model (i.e., in Figure 2) with weight , whose input is the hidden state from the last attention layer. The output is converted to a value between and through a sigmoid layer to represent the final prediction. We use L1 loss for HLP:
The full training objective is a weighted sum of NTP loss and HLP loss:
where is the tunable weight. In experiments, we set , which achieves good results across benchmarks empirically. We further study the relationship between the values of and the FIM performance of models trained with HLP. As discussed in Appendix A.1, the performance of HLP is robust to different values of .
Overhead Analysis. While HLP introduces the additional during training, the number of added parameters is < 0.01% of the base model, which incurs almost zero training overhead. Furthermore, the additional head will not be used during inference, leading to zero inference overhead.
4 Experiments
| SAFIM | Average | ||||
| Algo | Algo | Control | API | ||
| DS-1.3B | 39.8 | 42.4 | 52.4 | 56.1 | |
| + HLP | 41.3 | 46.1 | 53.4 | 59.0 | 50.0∗ |
| DS-6.7B | 53.4 | 56.7 | 66.6 | 69.0 | |
| + HLP | 53.5 | 57.4 | 66.9 | 69.7 | 61.9∗ |
| SC2-3B | 45.4 | 49.7 | 57.1 | 61.3 | |
| + HLP | 47.2 | 52.1 | 58.7 | 64.5 | 55.6∗ |
| SC2-7B | 48.4 | 53.1 | 60.4 | 63.9 | |
| + HLP | 49.4 | 54.5 | 61.8 | 65.8 | 57.9∗ |
Training. We conduct continual pre-training on a set of code LLMs of different model families and sizes to validate the effectiveness of HLP. Specifically, DeepSeek-Coder-Base 1.3B/6.7B (dscoder) and StarCoder2 3B/7B (starcodertwo) are involved in our experiments. We use AdamW (adamw) as the optimizer with and . We use a cosine learning rate scheduler with a peak learning rate equal to that at the pre-training end. All models are trained for 200K steps with a global batch size of 512.
Dataset. We use a subset of the-stack-v2 (starcodertwo) including Python, Java, C++, and C#. Following existing works (dscoder; starcodertwo), FIM rate is set to . We employ Best-fit Packing (ding2024fewertruncationsimprovelanguage) to group multiple files into each training sequence while masking out cross-file attention. The -- split is applied to each file independently rather than the whole training sequence.
| CrossCodeEval / CrossCodeLongEval | Average | |||||||
| Line | Chunk | Function | ||||||
| EM | ES | EM | ES | EM | ES | EM | ES | |
| DS-1.3B | 15.23 | 49.64 | 22.48 | 56.40 | 4.58 | 33.96 | ||
| + HLP | 18.99 | 55.47 | 24.32 | 58.77 | 5.12 | 35.25 | 16.14∗ | 49.83∗ |
| DS-6.7B | 26.23 | 62.07 | 28.90 | 62.37 | 7.50 | 41.42 | ||
| + HLP | 27.35 | 63.54 | 30.08 | 63.18 | 7.22 | 40.99 | 21.55∗ | 55.90∗ |
| SC2-3B | 24.17 | 59.89 | 22.20 | 52.69 | 6.80 | 38.13 | ||
| + HLP | 25.67 | 62.62 | 30.66 | 62.01 | 7.18 | 39.42 | 21.17∗ | 54.68∗ |
| SC2-7B | 26.00 | 61.68 | 27.14 | 56.52 | 7.66 | 39.54 | ||
| + HLP | 27.58 | 63.84 | 32.86 | 64.07 | 8.44 | 41.03 | 22.96∗ | 56.31∗ |
| RepoEval | Average | |||||||
| Line | API | Function | ||||||
| EM | ES | EM | ES | EM | ES | EM | ES | |
| DS-1.3B | 24.50 | 50.42 | 18.81 | 58.15 | 3.96 | 29.73 | ||
| + HLP | 27.25 | 53.45 | 21.81 | 59.79 | 5.93 | 31.92 | 18.33∗ | 48.39∗ |
| DS-6.7B | 26.62 | 52.59 | 22.69 | 61.94 | 7.47 | 36.24 | ||
| + HLP | 30.31 | 55.97 | 25.12 | 63.06 | 7.69 | 37.22 | 21.04∗ | 52.08∗ |
| SC2-3B | 21.88 | 46.74 | 18.81 | 56.66 | 4.40 | 29.99 | ||
| + HLP | 26.56 | 50.56 | 23.06 | 61.02 | 7.25 | 33.79 | 18.96∗ | 48.46∗ |
| SC2-7B | 27.94 | 51.60 | 21.56 | 58.98 | 6.81 | 32.80 | ||
| + HLP | 34.19 | 57.29 | 27.31 | 63.04 | 8.35 | 35.40 | 23.28∗ | 51.91∗ |
We conduct controlled experiments for all the studied code LLMs in our experiments. Specifically, we conduct two continual pre-training experiments for each model as follows:
NTP: existing pre-training scheme with next-token prediction (NTP) objective only.
NTP + HLP: our newly proposed objective that incorporates horizon-length prediction (HLP) objective with next-token prediction (NTP) objective.
Throughout this section, we determine the end of generation solely based on <eoi> generated by the model without any rule-based post-processing, unless otherwise specified (§4.4). We also conduct statistical analysis: results marked with * are statistically significant () based on paired t-tests.
4.1 Syntax-Aware Multilingual Code FIM
We use SAFIM (safim), a syntax-aware and multilingual code Fill-in-the-Middle benchmark, to evaluate the effectiveness of HLP. SAFIM focuses on syntax-aware completions of program structures, covering algorithmic block (i.e., Algo and Algo), control-flow expression (i.e., Control), and API function call (i.e., API). It consists of 17,720 examples from four different programming languages, including Python, Java, C++, and C#. SAFIM employs execution-based evaluation and reports pass@1 as the evaluation metric. As shown in Table 3, compared with NTP only, adding HLP achieves up to 5% improvements on average across all the studied code LLMs. Importantly, the improvement is consistent across languages and program structures, demonstrating HLP’s ability to enhance FIM capabilities regardless of language or completion context.
4.2 Repository-Level Cross-File Code FIM
In addition to single-file FIM evaluation with SAFIM, we also evaluate the effectiveness of HLP on repository-level code Fill-in-the-Middle in cross-file scenarios via CrossCodeEval (crosscodeeval), CrossCodeLongEval (repoformer), and RepoEval (repoeval). CrossCodeEval (Python) and CrossCodeLongEval are two repository-level cross-file benchmarks that leverage more than 1500 raw Python repositories to construct 12,665 examples across line, chunk, and function completion tasks, which are used for a more rigorous evaluation. RepoEval is another repository-level cross-file code completion benchmark consisting of 3,655 line, API, and function completion tasks created from 32 Python repositories. We follow existing work (repoformer) to evaluate the model’s FIM performance on these benchmarks and use Exact Match (EM) and Edit Similarity (ES) as our evaluation metrics. As shown in Table 4, adding HLP provides consistent improvements for all models across different benchmarks and completion tasks. Specifically, HLP achieves up to 24% improvements on EM and 9% improvements on ES relatively, showing its significant effectiveness.
| Code Repair | Code Reasoning | ||
| Defects4J | CRUXEval-I | CRUXEval-O | |
| DS-1.3B | 33 | ||
| + HLP | 39 | 44.7∗ | 31.8∗ |
| DS-6.7B | 58 | ||
| + HLP | 59 | 52.4 | 39.6 |
| SC2-3B | 39 | ||
| + HLP | 41 | 43.9∗ | 32.6∗ |
| SC2-7B | 41 | ||
| + HLP | 47 | 45.5∗ | 36.1∗ |
4.3 Code Repair via Fill-in-the-Middle
To assess HLP’s impact on practical applications beyond code completion, we evaluate the performance of code repair using Defects4J (defects). Defects4J consists of open-source bugs found across 15 Java repositories. Following existing works (fitrepair; chatrepair), we collect 313 single-hunk bugs from Defects4J that can be fixed by replacing or adding a continuous code hunk. Specifically, for each bug, models are prompted to generate the correct code hunk (i.e., patch) given the left and right contexts of the buggy code hunk, with correctness verified by project test suites. As shown in the “Code Repair” section of Table 5, adding HLP during training results in relatively up to 18% more bugs fixed by the model222Note that Defects4J is a small dataset with only 313 examples and models can only solve 30-60 out of those, which makes it hard to obtain statistically significant differences with greedy decoding., suggesting that enhanced planning capabilities translate directly to better bug-fixing performance.
4.4 Code Reasoning via Fill-in-the-Middle
We further examine whether HLP’s planning benefits extend to code reasoning tasks. We use CRUXEval (cruxeval) which comprises 800 Python functions paired with two distinct tasks: CRUXEval-I, where LLMs need to predict the input from the known output, and CRUXEval-O, where LLMs are required to predict the output based on the given input.
We reformat prompts of CRUXEval-I into FIM style and leave CRUXEval-O as L2R generation, both of which are evaluated in zero-shot setting. Different from previous subsections where post-processing is not used, we follow the same pipeline as in the original CRUXEval paper to extract accurate input/output values from generation because we are focusing on evaluating the reasoning capability of LLMs rather than their capability of generating correct code333In CRUXEval-I, we only want to evaluate the correctness of the input value infilled by LLMs in the given assertion. However, FIM-style prompts we use in the experiments does not restrict LLMs from writing multiple assertions before starting infilling the given assertion, which is useless in this task. So we use post-processing techniques to extract the input value infilled for the given assertion to better evaluate the reasoning capabilities.. As shown in the “Code Reasoning” section of Table 5, HLP demonstrates up to 6% improvements on both CRUXEval-I and CRUXEval-O tasks for all the code LLMs consistently, which shows that HLP also improves intrinsic code reasoning capabilities of LLMs.
5 Discussion
5.1 NTP Alone Cannot Yield Horizon Awareness
A key question of interest is whether standard next-token prediction (NTP) inherently provides models with awareness of generation horizons. Through systematic analysis, we demonstrate that hidden states of models trained with NTP alone do not capture meaningful information about required completion lengths.
To quantify this, we design a probing task that attempts to predict the remaining completion length from models’ hidden states. We extract hidden states from 20K code snippets (approximately 7.8M tokens) using models trained with and without HLP and split the data to ensure no sequence-level overlap between train and test. Using these hidden states as input and normalized remaining token counts as targets, we fit linear regression models while keeping the underlying transformer parameters frozen.
| Test | DS-1.3B | DS-6.7B | SC2-3B | SC2-7B |
| NTP | 0.440 | 0.519 | 0.356 | 0.410 |
| NTP+HLP | 0.915 | 0.913 | 0.932 | 0.932 |
Figure 3 shows predicted versus actual remaining token percentages at different positions, while Table 6 reports the coefficient of determination . Models trained with NTP alone show poor fit, indicating their hidden states lack horizon information. In contrast, models trained with HLP show much stronger correlations between hidden states and remaining lengths. This result demonstrates that horizon awareness must be explicitly trained rather than emerging naturally from NTP. We also carried out a non-linear probing experiment with the same inputs and targets by replacing linear regression models with an MLP regressor that uses logistic activation function (Appendix A.2), which further confirms our finding here.
5.2 Why Horizon-Length Prediction works?
The effectiveness of Horizon-Length Prediction (HLP) stems from its ability to address a critical limitation of standard next-token prediction (NTP) in autoregressive models: the absence of global planning awareness. While NTP optimizes for local token-level coherence by maximizing the likelihood of immediate next tokens, it inherently struggles to enforce long-horizon structural consistency between the generated and the given in FIM. This limitation becomes pronounced when the model must align generated code with both preceding logic and subsequent constraints (e.g., API calls or variable dependencies defined in the suffix).
HLP bridges this gap by teaching models to explicitly condition the model on the remaining generation horizon at every decoding step. By regressing the normalized count of future tokens required to complete , HLP effectively decomposes the infilling task into a sequence of length-aware subgoals, where each generated token is tied to a dynamically updated “budget” of remaining steps. This mechanism mirrors human problem-solving strategies, where progress estimation (e.g., "30% of steps remaining") guides iterative refinement of intermediate actions to meet global objectives.
The success of HLP is rooted in its complementary role to NTP. While NTP ensures local fluency, HLP provides a global scaffold for planning. By unifying local token prediction with global horizon awareness, HLP bridges the gap between autoregressive decoding and holistic reasoning, enabling models to dynamically adapt their generation strategy to long-horizon constraints. This synergy is particularly critical in code infilling, where the correctness of depends on both preceding logic (i.e., ) and subsequent context (i.e., ), demanding a balance between immediate token likelihood followed by and forward-looking structural coherence constrained by distant .
Our in-depth attention analysis provides direct evidence of this enhanced planning capability. Figure 4 shows that models trained with HLP pay significantly more attention to context, particularly at the start of the generation. This indicates that HLP enables proactive planning by considering long-horizon constraints from the beginning. Furthermore, the consistent improvements across FIM, bug-fixing, and reasoning tasks demonstrate that this enhanced planning capability generalizes beyond simple completion scenarios.
In summary, HLP transforms code generation from a myopic token-by-token process into a structured planning task. It enables models to dynamically adapt their generation strategy to long-horizon constraints, ensuring that each step not only satisfies local fluency but also contributes to a globally coherent solution. This dual focus on immediate actions and overarching goals positions HLP as a pivotal advancement in code LLMs.
5.3 HLP Mitigates Planning Failures in FIM
We conduct a quantitative analysis to empirically demonstrate the frequency of planning failures in FIM and evaluate the effectiveness of HLP in addressing these failures. As presented in Figure 1 (§2), a prevalent manifestation of planning failures in FIM is the "correct code but incorrect order" pattern, wherein the model generates the final line of ground truth code accurately but positions it incorrectly within the completion. To maintain analytical tractability, our study focuses specifically on planning failures that result in this particular pattern. We establish the following criteria to identify instances of the "correct code but incorrect order" pattern in model’s FIM completions:
Completion fails to pass all the unit tests.
Completion contains the last line from ground truth.
Completion does not end with this last line.
Testing on SAFIM’s Algorithmic and Algorithmic_v2 partitions which focus on multi-line FIM, we find that models trained without HLP exhibit the “correct code but incorrect order” pattern in 5-10% of problems across the benchmark, as shown in Table 7. With HLP training, this drops to 4-6%, demonstrating that HLP helps alleviate lack-of-planning failures effectively.
| SAFIM | Avg | ||
| Algo | Algo | ||
| DS-1.3B | 9.34% | 10.48% | 9.91% |
| + HLP | 6.83% | 6.52% | 6.68% |
| DS-6.7B | 5.16% | 5.93% | 5.55% |
| + HLP | 4.36% | 4.44% | 4.40% |
| SC2-3B | 7.25% | 7.35% | 7.30% |
| + HLP | 5.65% | 5.43% | 5.54% |
| SC2-7B | 5.79% | 6.26% | 6.03% |
| + HLP | 4.69% | 4.29% | 4.49% |
6 Related Work
Fill-in-the-Middle for Code Language Models Large language models trained on massive source code data have demonstrated great potential in various applications for software development. While early models such as Codex (humaneval) and CodeGen (codegen) only support Left-to-Right (L2R) generation, Fill-in-the-Middle (or infilling) has attracted increased attention because right context naturally carries an indispensable part of information for completing code in the middle (incoder; openaifim). Subsequently, FIM training has become a common practice widely adopted by most code LLMs, such as StarCoder (starcoder; starcodertwo), DeepSeek-Coder (dscoder; dscodervtwo), and Code Llama (codellama).
Existing models generally tackle the infilling problem by breaking a code snippet into --, and reordering them into -- (PSM) or -- (SPM), which are then used for next-token prediction (NTP) training. We point out that the infilling task cannot be effectively learned with NTP alone, as it requires planning capability for the model to fluently and meaningfully connect the middle completion to through forward looking during auto-regressive decoding.
An alternative approach is to train two language models in different directions, with one generating from left to right and the other from right to left, and have the two generations meet in the middle (meet-in-the-middle). Nevertheless, the L2R model does not have access to the right context, and vice versa, which impedes holistic planning that takes into account the context from both sides.
Planning and Lookahead in Language Generation Standard decoder-only models are trained with next-token prediction and used to sequentially predict one token at a time, conditioned only on past tokens, in an auto-regressive manner. One drawback of this paradigm is that models are not aware of future tokens during decoding. The token that maximizes the conditional probability at current step may lead to suboptimal continuation, and consequently the model can fail to compose a fluent and sensible generation that meets human requirements. Various decoding techniques have been proposed to address the problem through tree search with lookahead heuristics, particularly for constrained generation problems (NeuroLogic; DeAL). While these methods are training-free, they inevitably incur additional cost of inference complexity.
Apart from those, multitoken proposed to predict multiple tokens from a single hidden state during both training and inference, which was shown to achieve stronger performance on coding tasks with no computation overhead. While multi-token prediction enhances models’ planning capability within the tokens predicted together (), we argue that with a small , the limited horizon is usually insufficient for planning in the case of infilling as the connection from to only happens towards the end of the generation. In contrast, HLP adopts a global and arbitrary long horizon over all future tokens by counting the remaining generation budget, which more effectively helps models to close the generation fluently with early planning444Please refer to Appendix §A.3 for more details..
7 Conclusion and Future Work
Fill-in-the-Middle is ubiquitous in code completion, reflecting the iterative nature of software development where code is frequently inserted between existing sections, and thus has become an important consideration in the development of code language models. The current FIM training paradigm splits and reorders training sequences for next-token prediction. However, this approach frequently results in models struggling to generate content that smoothly aligns with the right context. While existing FIM benchmarks rely on different post-processing methods to circumvent this problem, we emphasize that such methods typically require dataset-specific assumptions, which are impractical in real-world scenarios.
To address this limitation and enhance the infilling capability of code language models, we propose Horizon-Length Prediction (HLP). HLP teaches models to predict the portion of remaining tokens at every step, enabling them to develop planning awareness over arbitrarily long horizons. Experiments across different model families and sizes show that HLP improves infilling performance on diverse FIM benchmarks, across file-level and repository-level, and without using any dataset-specific post-processing. Moreover, the enhanced planning capability acquired through HLP training also boosts models’ performance on code reasoning tasks, suggesting broader benefits for language models’ reasoning capabilities. Importantly, HLP achieves these improvements while maintaining efficiency, with negligible training overhead and no additional inference cost.
Our work marks a significant advancement in developing more effective code language models for real-world applications. Future work could explore extending HLP to other generation tasks requiring long-horizon planning and investigate its potential for enhancing general reasoning capabilities.
Limitations
While HLP has proven effective through extensive evaluation in the paper, our pre-training experiments are restricted to models of 7B parameters due to the computation budget. It is prohibitively expensive to perform pretraining experiments for LLM, and unfortunately we do not have enough resources to demonstrate the impact of HLP on larger models. This highlights a broader challenge in open science for pre-training research. In addition, this work has mainly focused on the code domain, but the idea of horizon-length prediction can be broadly applicable for improving models’ reasoning capability in natural language and mathematical domains as well, which we leave to future work.
Appendix A Appendix
A.1 Effect of Value Selection
Following the original experimental settings (but with 100k training steps given limited computational resources), we studied a series of values and report both HLP validation loss and SAFIM pass@1. As shown in Table 8, our method is robust to hyperparameter selection: all non-zero values improve over the baseline (). Moreover, we observe that HLP loss quickly decreases within the first few thousand training steps with all positive ’s, indicating that a large isn’t necessary for models to learn HLP well. This also coincides with our observation on downstream performance: while lower HLP loss generally correlates with better SAFIM performance, the improvement plateaus beyond a certain threshold, suggesting that once the model acquires basic horizon awareness, further optimization of HLP loss provides diminishing returns for downstream performance.
| HLP Valid Loss | SAFIM (pass@1) | |
| 0 | 0.254 | 47.40% |
| 0.01 | 0.060 | 48.98% |
| 0.1 | 0.048 | 49.10% |
| 0.2 | 0.045 | 48.74% |
| 0.5 | 0.043 | 49.34% |
A.2 Non-Linear Probing for Horizon Awareness
By replacing the linear regression models in the original linear probing experiment with an MLP regressor that uses logistic activation function, we further conduct a non-linear probing experiment to study whether NTP is able to yield horizon awareness. As shown in Table 9, even under the assumption of the non-linear relationship, models trained without HLP cannot effectively extract horizon information from their hidden states. This demonstrates that horizon awareness must be explicitly trained rather than emerging naturally from NTP.
| Test | DS-1.3B | DS-6.7B | SC2-3B | SC2-7B |
| NTP | 0.392 | 0.431 | 0.180 | 0.226 |
| NTP+HLP | 0.897 | 0.885 | 0.856 | 0.871 |
A.3 Comparing HLP with Multi-token Prediction
| SAFIM | Average | ||||
| Algo | Algo | Control | API | ||
| DS-1.3B | 39.8 | 42.4 | 52.4 | 56.1 | 47.7 |
| + HLP | 41.3 | 46.1 | 53.4 | 59.0 | 50.0 |
| + multi-token prediction | 38.1 | 41.3 | 51.7 | 55.8 | 46.7 |
As discussed in §6, we argue that multi-token prediction (multitoken) is insufficient for planning in FIM, because multi-token prediction only enhances models’ planning capability for a short and limited horizon, which does not suites FIM well as the connection from to happens over a long horizon. Instead, HLP focuses on long-horizon planning and is more effective for FIM. We conduct an experiment following the same settings in §4 to compare HLP with multi-token prediction on DeepSeek-Coder-Base 1.3B by predicting the next 4 tokens (multitoken). We report their performance on SAFIM. As shown in Table 10, while adding HLP to NTP largely improves the model’s performance on SAFIM, multi-token prediction fails to do so. These results provide empirical evidence that long-horizon planning capabilities brought by HLP is essential for advancing FIM performance.
A.4 HLP Improves Self-Infilling Performance
Recently, some strategies have been proposed to conduct L2R code generation based on FIM capabilities and self-infilling (zheng2024selfinfillingcodegeneration) is one of the representative strategies of this kind. Given an initial input prompt , self-infilling performs L2R code generation by using FIM decoding to (1) generate based on and then (2) generate based on and previously generated . Furthermore, self-infilling proposes a looping mechanism to improve the generated code iteratively, where it first uses L2R decoding to generate a new ′ based on and previously generated and then uses FIM decoding to generate a new ′ based on and this newly generated ′. This looping procedure can be continued for multiple rounds to obtain greater improvements.
| HumanEval | |||||
| DS-1.3B | 27.4 | 29.3 | 31.7 | 31.7 | 31.7 |
| + HLP | 30.5 | 31.7 | 31.7 | 32.3 | 32.9 |
Since self-infilling relies on FIM capabilities for L2R code generation, it is interesting to study whether HLP can also enhance its performance. To this end, we evaluate the self-infilling performance of DeepSeek-Coder-Base 1.3B trained with and without HLP on HumanEval using greedy decoding, following the setting of the original paper (zheng2024selfinfillingcodegeneration). We report models’ performance on HumanEval with ranging from 0 to 4, where denotes the number of times the decoding process goes through the loop and represents that the looping mechanism is not activated. As shown in Table 11, the model trained with HLP consistently outperforms the model trained without HLP across different number of self-improving steps. Furthermore, with more self-improving steps, while the performance of the model trained without HLP gets stuck, the performance of the model trained with HLP continues to showcase steady improvements.
A.5 Effect of HLP on Left-to-Right Performance
While HLP have significantly improved the FIM performance of LLMs, we also study its impact on the L2R code completion. To this end, we evaluate L2R performance on HumanEval (humaneval) and MBPP (mbpp) with DeepSeek-Coder-Base 1.3B. We further employ HumanEval+ and MBPP+ from EvalPlus evalplus for more rigorous evaluation with better test coverage. As shown in Table 12, with HLP applied to FIM data only (i.e., ), the performance on L2R tasks sometimes shows a slight degradation. We hypothesize that applying HLP to only causes unbalanced training on and parts.
| Left-to-Right | Fill-in-the-Middle | ||
| HumanEval (+) | MBPP (+) | SAFIM | |
| DS-1.3B | 26.3 (22.0) | 45.8 (36.7) | 47.7 |
| + HLP | 25.5 (21.3) | 45.8 (36.5) | 50.0 |
| + HLP + HLP | 26.2 (22.0) | 45.7 (36.6) | 49.6 |
| SAFIM | Average | ||||
| Algo | Algo | Control | API | ||
| DS-1.3B | 39.8 | 42.4 | 52.4 | 56.1 | 47.7 |
| + HLP (linear) | 41.3 | 46.1 | 53.4 | 59.0 | 50.0 |
| + HLP (mlp) | 41.6 | 45.9 | 54.0 | 57.4 | 49.7 |
| SAFIM | Average | ||||
| Algo | Algo | Control | API | ||
| DS-1.3B | 39.8 | 42.4 | 52.4 | 56.1 | 47.7 |
| + HLP (all) | 41.3 | 46.1 | 53.4 | 59.0 | 50.0 |
| + HLP (first) | 40.0 | 44.4 | 52.3 | 57.4 | 48.5 |
| SAFIM | Average | ||||
| Algo | Algo | Control | API | ||
| DS-1.3B | 39.8 | 42.4 | 52.4 | 56.1 | 47.7 |
| + HLP (normalized) | 41.3 | 46.1 | 53.4 | 59.0 | 50.0 |
| + HLP (raw) | 27.7 | 29.8 | 34.4 | 47.7 | 34.9 |
To mitigate such effect, we need to devise another HLP task that can be applied to L2R training (i.e., ). However, the original design of HLP in §3 is not directly applicable to L2R data. While the end of in FIM data is strictly bounded by the beginning of , the end of L2R data does not have any clear signals, as it is often possible to add additional contents (e.g., another line of code or a new helper function) to the end of document fluently without any restrictions.
Therefore, instead of taking the entire code file as the prediction horizon, we ask the model to predict the number of future tokens required to complete current line in L2R training, which is a natural semantic unit in code. Furthermore, to avoid conflicts between and , we use two independent s to let the model learn and separately. As shown in Table 12, by applying and simultaneously, the performance degradation on L2R tasks is recovered, with the improvement on FIM tasks largely retained. These results demonstrate the generalizable effectiveness of HLP and shows the huge potential of applying the idea of HLP to more general training scenarios.
A.6 Additional Ablation Studies
We conduct several ablation studies to justify the design choices of HLP. In this section, we conduct experiments using DeepSeek-Coder-Base 1.3B, follow the same settings in §4, and report models’ performance on SAFIM.
Complexity of . We conduct an experiment to study the effect of the complexity of by replacing the original linear layer (i.e., “HLP (linear)”) with a two-layer MLP with ReLU (i.e., “HLP (mlp)”). As shown in Table 13, increasing the complexity of does not bring significant improvements. We have also conducted a paired t-test between “HLP (linear)” and “HLP (mlp)” and did not see any clear directional statistical significance between them. Such results indicate that the complexity of does not have a major impact on performance.
Applying HLP to all tokens v.s. first token only. While it is easy to see that knowing the HLP loss on the first token is sufficient to infer the horizon length in theory, having HLP loss on every token provides denser and more consistent supervision signals which makes learning easier (as discussed in §3). It also helps regularize the hidden representation of every subsequent token. To empirically show this, we conduct an experiment by applying HLP to the first token only (i.e., “HLP (first)”) and compared its performance with our original HLP design (i.e., “HLP (all)”). As shown in Table 14, while applying HLP only to the first token performs better than NTP only, applying HLP loss for each token can achieve better performance than applying it to just the first token.
Normalized v.s. unnormalized targets. We use normalized targets in our original HLP design (i.e., using rather than ) is that the scale of HLP loss will be otherwise in a huge range, e.g., some examples has single digit loss while some might have thousands. To further study the effect of normalization, we conduct an experiment by using as the target (i.e., “HLP (raw)”) rather than (i.e., “HLP (normalized)”). To achieve this, we remove the sigmoid function from the original HLP. As shown in Table 15, setting the target as fails to improve FIM performance, likely due to the large-scale HLP loss after using as the target interferes with NTP pre-training.